OpenStudy (anonymous):

Suppose that $2000 is loaned at a rate of 7%, compounded quarterly. Assuming that no payments are made, find the amount owed after 7 years. Do not round any intermediate computations, and round your answer to the nearest cent.

Parth (parthkohli):

All right. 7 years is supposed to be \(7 \times 4 = 28\) quarters.

Parth (parthkohli):

Now, \(7\% = 0.7\) We can now use the continuous compounding formula.

Parth (parthkohli):

\( \color{Black}{\Rightarrow 2000(1 + 0.07)^{28}}\)

Parth (parthkohli):

I meant \(7\% = 0.07\).

Parth (parthkohli):

Use your calculator to solve \(2000 \times 1.07^{28}\).

OpenStudy (anonymous):

Would it be 13297.6767 ?

OpenStudy (anonymous):

There are 4 quarters in a year i.e 3 months is one quarter (1/4 of a year).

OpenStudy (anonymous):

2000*(1+.07/4)^28

OpenStudy (anonymous):

3250.83

OpenStudy (anonymous):

4*3*7=84% 2000*84%=3680